OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3 π (x - 7)2 (x - 3) What does the factor 1 over 3 (x - 7) represent? (4 points) The total surface area of the tank One third of the radius of the tank One third of the height of the tank The total area of the base of the tank

OpenStudy (anonymous):

r=h-4 if h = x-3 then r= x-7 and V= 1/3 r.h now try

OpenStudy (anonymous):

volume =1/3*pi*r^2*h V= (1/3*h) (pi*r^2) V=1/3*(x-3) (pi*(x-7)^2)

OpenStudy (anonymous):

1/3(x-7)=1/3 r

OpenStudy (anonymous):

One third of the radius of the tank